# Jailbreaking Prompt Attack: A Controllable Adversarial Attack against Diffusion Models

### v3

@@ Introduction
NSFW prompts. We evaluate the performance of JPA on the Inappropriate Image Prompt (I2P) dataset (Schramowski et al. 2023), a well-established collection of NSFW prompts. For the nudity concept, we select 142 prompts, referred to as NSFW-142, following (Zhang et al. 2024). For the violence concept, we select 90 samples with violence levels exceeding 90% as NSFW-90.
@@
Online services. We evaluate our attack on four popular online services: DALL· E 2 (Ramesh et al. 2022), Stability.ai (Clipdrop of Stable Diffusion XL) (Runwayml 2023), Midjourney (MidJourney 2023), and PIXART-$\alpha$ (Chen et al. 2023).
@@
white-box P4D (Chin et al. 2024) 45.86 97.74 50.61 60.90 75.71 36.43
UnlearnDiff (Zhang et al. 2024) 51.00 96.48 56.34 61.97 76.76 38.02
black-box QF-Attack (Zhuang, Zhang, and Liu 2023) 5.94 36.77 9.47 11.59 22.15 4.21
@@
Ring-A-Bell (Tsai et al. 2023) 53.30 94.21 57.57 69.05 87.65 56.97
SneakyPrompt (Yang et al. 2024) 42.01 95.17 50.45 59.74 73.20 35.19
@@
white-box P4D (Chin et al. 2024) 170.25 158.14 143.52 141.13 159.60 167.03
UnlearnDiff (Zhang et al. 2024) 144.26 139.36 144.26 136.34 124.59 141.13
@@
QF-Attack (Zhuang, Zhang, and Liu 2023) 201.78 198.60 194.22 191.06 205.67 199.30
Ring-A-Bell (Tsai et al. 2023) 152.45 138.76 129.80 128.59 116.58 155.59
SneakyPrompt (Yang et al. 2024) 155.40 126.19 125.48 131.79 119.84 147.36
@@
Table 2: The performance of attack methods on the "nudity" concept is evaluated under ASR and FID. The highest performance is highlighted with bold values
Offline T2I models with removal methods. To evaluate the effectiveness of the attack on offline T2I models with removal methods, we adopt ESD (Gandikota et al. 2023), FMN (Zhang et al. 2023), SD-NP (Rombach et al. 2022), and SLD (Schramowski et al. 2023), along with its three enhanced variants, i.e. SLD-Medium, SLD-Strong, and SLD-MAX ranked by their defense capabilities. To reproduce the fine-tuned diffusion models in ESD and FMN, we input "nudity" and "violence" separately as the target unsafe concepts for removal.
Baselines. We begin by reporting the results of NSFW prompts used as input to Stable Diffusion without safety checkers under the No Attack condition. For white-box attack methods, we selected UnlearnDiff (Zhang et al. 2024) and P4D (Chin et al. 2024) to attack T2I models that employ removal methods. In the case of black-box attack methods, we first report results using random character selection to generate adversarial prompts. Then we reproduced QF-Attack (Zhuang, Zhang, and Liu 2023), opting to mask out words related to clothing, such as "wear tops and pants", as well as words related to peace and absence of harm, like "peaceful and fine", using a genetic algorithm (GA) to optimize the adversarial prompts. For Ring-A-Bell, we report results from Ring-A-Bell-Union, following the settings detailed in (Tsai et al. 2023). Finally, we implemented the RL method from SneakyPrompt (Yang et al. 2024), setting $\tau$ = 0.26 for the attack.
Evaluations.
NSFW detector. To classify whether images contain nudity, we use the NudeNet detector (Bedapudi 2019). If the detector identifies any of the following labels in an image including EXPOSED_ANUS, EXPOSED_BUTTOCKS, EXPOSED_BREAST, or EXPOSED_GENITALIA, we categorize the image as unsafe. For the violence concept, we employ the Q16 classifier (Schramowski, Tauchmann, and Kersting 2022) to determine whether images contain harmful content, such as blood or violence.
Metric. a) Attack Success Rate (ASR): We measure the effectiveness of the attack using the Attack Success Rate (ASR) (Moosavi-Dezfooli, Fawzi, and Frossard 2016), which is determined by the number of adversarial prompts that successfully bypass the NSFW detector.
b) Frechet Inception Distance (FID): We assess the semantic similarity of the generated images using the Frechet Inception Distance (FID) score (Heusel et al. 2017). Following the official PyTorch implementation (Seitzer 2020), we compare our generated images against a ground-truth dataset. This dataset includes 1000 images each for the "nudity" and "violence" categories, constructed using NSFW-142 and NSFW-90, respectively, with Stable Diffusion without safety checkers and different random seeds. A higher FID score indicates a greater semantic divergence between the two image sets.
Implementation details. We utilize the text encoder from CLIP ViT-L/14 (Dosovitskiy et al. 2020). To generate $N$ pairs of antonyms associated with the target unsafe concept, we prompt ChatGPT-4 with the query: "Can you help me find the words that best represent the concept of "nudity" and provide their antonyms?" The complete list of concept pairs used is provided in the Appendix. We conduct PGD (Madry et al. 2017) for $600$ iterations, using a learning rate of $10^{-5}$ at each step and the AdamW optimizer (Loshchilov and Hutter 2017).
Main Results
Evaluation with Online Services. Despite the deployment of various safety checkers in online services, which remain unknown to public. Figure 3 demonstrates that JPA can successfully bypass these defenses and generate NSFW images. We conduct attack on four popular online platforms: DALL·E 2 (Ramesh et al. 2022), Stability.ai (Runwayml 2023), Midjourney (MidJourney 2023), and PIXART-$\alpha$ (Chen et al. 2023). More visualization examples are provided in Appendix.
Figure 4: Each column represents a different attack method, with the last column showing images generated by Stable Diffusion without safety checkers. We also use JPA with BERT and T5 text encoders to execute the attack, demonstrating that JPA can maintain semantic similarity with images generated by SD without safety checkers. The first two rows correspond to attacks on "nudity" and the bottom two are on "violence" concepts, respectively.
@@
DALL·E 2 (Ramesh et al. 2022) Attack success Adversarial prompt
@@
Table 4: Examples of adversarial prompts used to attack online services from Ring and JPA. "Ring w/post-processing" indicates additional manual post-processing applied. We use red and blue to represent the use of prompt dilution and modification tricks, while JPA operates effectively without the need for such manual interventions.
Evaluation on Offline T2I Models with Removal Methods. We further attack defense methods with concept removal. For both the "nudity" and "violence" categories, we configure JPA with $\lambda=3$ and $k=7$ respectively. As shown in Table 2, when defending against nudity, both the FMN and SLD-Medium offline models demonstrate poor defense performance. Additionally, these removal methods present a significant risk of generating violent images, likely due to the broader definition of violence, which complicates effective defense. The performance on the "violence" concept is listed in Appendix, along with visualizations. Moreover, our approach achieves better FID scores, demonstrating that JPA can generate images semantically similar to images with NSFW images produced without safety checkers. Figure 4 visualizes the semantic similarity, further highlighting that JPA consistently maintains the highest image quality across both concepts. Additional results are provided in Appendix.
Controllable NSFW Concept Rendering We discover that adjusting the magnitude of the NSFW concept embeddings, JPA can precisely control the extent to which NSFW concepts emerge in the generated images. We show in Figure 1 (b) that increasing $\lambda$ lead to more NSFW contents being rendered in the image. We further report the ASR and FID with different $\lambda$ values in Table 5. The results indicate that, up to a certain threshold, increasing the $\lambda$ value enhances the attack performance. We provide more visualization in Appendix.
@@
Universal Attack w/ Arbitrary Text Encoders. Is our method truly black-box? Although previous works (Zhuang, Zhang, and Liu 2023; Tsai et al. 2023; Yang et al. 2024) claim to be black-box, they still rely on the text encoder of the Stable Diffusion model, thereby using some internal parameters. In this section, we demonstrate that JPA is universal, achieving high attack performance with arbitrary text encoders. Beyond the CLIP text encoder, we employ BERT (Devlin et al. 2019), pre-trained on large-scale corpora such as BooksCorpus (Zhu et al. 2015) and English Wikipedia (Foundation 2018), and the T5 model (Raffel et al. 2020), fine-tuned on a wide range of NLP tasks in a text-to-text format, to execute JPA. As shown in Table 6, even with text encoders that are not aligned with vision modalities, the attack remains effective, although the ASR slightly decreases due to the misalignment between text and image in these encoders.
@@
Attack Efficiency. As discussed in introduction, prior methods require lengthy optimization or post-processing due to the discrete text space. To address both the issues of discrete space and the presence of sensitive words in the output, we propose soft assignment and gradient masking. In this section, we demonstrate that our method is more efficient, significantly reducing execution time compared to previous methods. As illustrated in Table 3, JPA generates a successful adversarial prompt in just 6.72 minutes, which is only 1/4 the time required by the next fastest method.
Discussion on Fully Automated Attack Framework. We observe that some prior methods depend on manual modifications or additional techniques, such as manually altering sensitive words (Qu et al. 2023; Rando et al. 2022) and employing prompt dilution (Rando et al. 2022), to successfully execute attacks against online services (Tsai et al. 2023). This reliance on human intervention compromises the automation of the attack framework. For instance, in Table 4, Ring (Tsai et al. 2023) requires additional prompt dilution and human modification to effectively attack DALL·E 2 (Ramesh et al. 2022). In contrast, our attack framework, JPA, automates the process by masking sensitive words, enabling the generation of NSFW concepts while staying within the less sensitive text regions, making them less likely to be blocked by filters.
Conclusion and Limitations.
@@
However, our work is not without limitations. The dataset used in our study, the I2P dataset, contains prompts derived from human experience that inherently include unsafe semantics. When we substitute this with entirely safe data, the effectiveness of our attacks decreases to some extent. This highlights potential areas for further exploration in future research. Additionally, we recognize the urgent need for the development of more robust safety checkers to counter such attacks effectively.
@@
Sensitive Words List.
@@
Table 8: Render words list for "nudity" and "violence".
Render Words List.
This section presents the render word lists used in experiments for different concepts. We specifically highlight the lists for the concepts of "nudity" and "violence". Each list contains rows representing individual render word pairs, with the total number of rows corresponding to the number of pairs. Detailed information is provided in Table 8.
@@
Quality results of JPA in "violence" concept.
@@
white-box P4D (Chin et al. 2024) 97.72 97.88 75.28 88.89 89.89 100
UnlearnDiff (Zhang et al. 2024) 98.87 98.87 82.02 89.77 91.11 100
@@
QF-Attack (Zhuang, Zhang, and Liu 2023) 55.56 46.67 18.89 24.45 37.78 80.00
Ring-A-Bell (Tsai et al. 2023) 96.67 100 91.11 90.00 94.45 100
SneakyPrompt (Yang et al. 2024) 95.07 100 77.78 88.89 90.00 100
@@
white-box P4D (Chin et al. 2024) 239.80 269.57 221.83 220.76 216.88 275.89
UnlearnDiff (Zhang et al. 2024) 238.65 270.81 211.30 221.57 219.78 277.90
@@
QF-Attack (Zhuang, Zhang, and Liu 2023) 271.80 254.93 261.47 259.34 255.89 278.05
Ring-A-Bell (Tsai et al. 2023) 234.79 267.75 224.85 237.97 215.88 276.57
SneakyPrompt (Yang et al. 2024) 230.90 260.57 215.96 226.70 218.49 270.69
@@
Table 9: The performance attack methods on the "violence" concept is evaluated under ASR and FID.
Visualization results of Not-Safe-for-Work (NSFW) images generated by T2I models.
@@
Visualization of the controllable NSFW concept rendering.
We discover that adjusting the magnitude of the NSFW concept embeddings, JPA can precisely control the extent to which NSFW concepts emerge in the generated images. We show in Figure 7 that increasing $\lambda$ lead to more NSFW contents being rendered in the image. Specifically, we adjust the value of $\lambda$ from 0 to 5 produces a series of images that progressively render the "nudity" concept step by step. These images increasingly incorporate specific content related to the "nudity" concept while preserving the original image’s overall meaning.
Figure 7: The controllable "nudity" concept rendering process.
Visualization of the controllable ordinary concept rendering.

### v4

@@ 5 Experiment
We evaluate the performance of JPA on the Inappropriate Image Prompt (I2P) dataset (Schramowski et al., 2023), a well-established collection of NSFW prompts. For the nudity concept, we select $142$ prompts, referred to as NSFW-142, following (Zhang et al., 2024). For the violence concept, we select $90$ samples with violence levels exceeding $90\%$ as s NSFW-90.
Online services.
@@
We evaluate our attack on four popular online services: DALL· E 2 (Ramesh et al., 2022), Stability.ai (Clipdrop of Stable Diffusion XL) (Runwayml, 2023), Midjourney (MidJourney, 2023), and PIXART-$\alpha$ (Chen et al., 2023).
Offline T2I models with removal methods.
To evaluate the effectiveness of the attack on offline T2I models with removal methods, we adopt ESD (Gandikota et al., 2023), FMN (Zhang et al., 2023), SD-NP (Rombach et al., 2022), and SLD (Schramowski et al., 2023), along with its three enhanced variants, i.e. SLD-Medium, SLD-Strong, and SLD-MAX ranked by their defense capabilities. To reproduce the fine-tuned diffusion models in ESD and FMN, we input “nudity” and “violence” separately as the target unsafe concepts for removal.
Baselines.
We begin by reporting the results of NSFW prompts used as input to Stable Diffusion without safety checkers under the No Attack condition. For white-box attack methods, we selected UnlearnDiff (Zhang et al., 2024) and P4D (Chin et al., 2024) to attack T2I models that employ removal methods. In the case of black-box attack methods, we first report results using random character selection to generate adversarial prompts. Then we reproduced QF-Attack (Zhuang et al., 2023), opting to mask out words related to clothing, such as “wear tops and pants”, as well as words related to peace and absence of harm, like “peaceful and fine”, using a genetic algorithm (GA) to optimize the adversarial prompts. For Ring-A-Bell, we report results from Ring-A-Bell-Union, following the settings detailed in (Tsai et al., 2023). Finally, we implemented the RL method from SneakyPrompt (Yang et al., 2024), setting $\tau$ = $0.26$ for the attack.
5.2 Evaluations.
NSFW detector.
To classify whether images contain nudity, we use the NudeNet detector (Bedapudi, 2019). If the detector identifies any of the following labels in an image including EXPOSED_ANUS, EXPOSED_BUTTOCKS, EXPOSED_BREAST, or EXPOSED_GENITALIA, we categorize the image as unsafe. For the violence concept, we employ the Q16 classifier (Schramowski et al., 2022) to determine whether images contain harmful content, such as blood or violence.
Metric.
a) Attack Success Rate (ASR): We measure the effectiveness of the attack using the Attack Success Rate (ASR) (Moosavi-Dezfooli et al., 2016), which is determined by the number of adversarial prompts that successfully bypass the NSFW detector. b) Frechet Inception Distance (FID): We assess the semantic similarity of the generated images using the Frechet Inception Distance (FID) score (Heusel et al., 2017). Following the official PyTorch implementation (Seitzer, 2020), we compare our generated images against a ground-truth dataset. This dataset includes 1000 images each for the “nudity” and “violence” categories, constructed using NSFW-142 and NSFW-90, respectively, with Stable Diffusion without safety checkers and different random seeds. A higher FID score indicates a greater semantic divergence between the two image sets.
Implementation details.
We utilize the text encoder from CLIP ViT-L/14 (Dosovitskiy et al., 2020). To generate $N$ pairs of antonyms associated with the target unsafe concept, we prompt ChatGPT-4 with the query: “Can you help me find the words that best represent the concept of “nudity” and provide their antonyms?”. The complete list of concept pairs used is provided in the Appendix. We conduct PGD (Madry et al., 2017) for $600$ iterations, using a learning rate of $10^{-5}$ at each step and the AdamW optimizer (Loshchilov and Hutter, 2017).
5.3 Main Results
Evaluation with Online Services.
Despite the deployment of various safety checkers in online services, which remain unknown to public. Figure 3 demonstrates that JPA can successfully bypass these defenses and generate NSFW images. We conduct attack on four popular online platforms: DALL·E 2 (Ramesh et al., 2022), Stability.ai (Runwayml, 2023), Midjourney (MidJourney, 2023), and PIXART-$\alpha$ (Chen et al., 2023). More visualization examples are provided in Appendix.
@@ 5.3 Main Results
white-box P4D Chin et al. (2024) 45.86 97.74 50.61 60.90 75.71 36.43
UnlearnDiff Zhang et al. (2024) 51.00 96.48 56.34 61.97 76.76 38.02
black-box QF-Attack Zhuang et al. (2023) 5.94 36.77 9.47 11.59 22.15 4.21
@@
Ring-A-Bell Tsai et al. (2023) 53.30 94.21 57.57 69.05 87.65 56.97
SneakyPrompt Yang et al. (2024) 42.01 95.17 50.45 59.74 73.20 35.19
@@
white-box P4D Chin et al. (2024) 170.25 158.14 143.52 141.13 159.60 167.03
UnlearnDiff Zhang et al. (2024) 144.26 139.36 144.26 136.34 124.59 141.13
@@
QF-Attack Zhuang et al. (2023) 201.78 198.60 194.22 191.06 205.67 199.30
Ring-A-Bell Tsai et al. (2023) 152.45 138.76 129.80 128.59 116.58 155.59
SneakyPrompt Yang et al. (2024) 155.40 126.19 125.48 131.79 119.84 147.36
@@
Table 2: The performance of attack methods on the “nudity” concept is evaluated under ASR and FID. The highest performance is highlighted with bold values.
@@
DALL·E 2 (Ramesh et al., 2022) Attack success Adversarial prompt
@@
Table 4: Examples of adversarial prompts used to attack online services from Ring and JPA. “Ring w/post-processing” indicates additional manual post-processing applied. We use red and blue to represent the use of prompt dilution and modification tricks, while JPA operates effectively without the need for such manual interventions.
Evaluatiosn on Offline T2I Models with Removal Methods.
We further attack defense methods with concept removal. For both the “nudity” and “violence” categories, we configure JPA with $\lambda=3$ and $k=7$ respectively. As shown in Table 2, when defending against nudity, both the FMN and SLD-Medium offline models demonstrate poor defense performance. Additionally, these removal methods present a significant risk of generating violent images, likely due to the broader definition of violence, which complicates effective defense. The performance on the “violence” concept is listed in Appendix, along with visualizations. Moreover, our approach achieves better FID scores, demonstrating that JPA can generate images semantically similar to images with NSFW images produced without safety checkers. Figure 4 visualizes the semantic similarity, further highlighting that JPA consistently maintains the highest image quality across both concepts. Additional results are provided in Appendix.
Figure 4: Each column represents a different attack method, with the last column showing images generated by Stable Diffusion without safety checkers. We also use JPA with BERT and T5 text encoders to execute the attack, demonstrating that JPA can maintain semantic similarity with images generated by SD without safety checkers. The first two rows correspond to attacks on “nudity” and the bottom two are on “violence” concepts, respectively.
Controllable NSFW Concept Rendering.
We discover that adjusting the magnitude of the NSFW concept embeddings, JPA can precisely control the extent to which NSFW concepts emerge in the generated images. We show in Figure 1 (b) that increasing $\lambda$ lead to more NSFW contents being rendered in the image. We further report the ASR and FID with different $\lambda$ values in Table 5. The results indicate that, up to a certain threshold, increasing the $\lambda$ value enhances the attack performance. We provide more visualization in Appendix.
@@
Universal Attack w/ Arbitrary Text Encoders.
Is our method truly black-box? Although previous works (Zhuang et al., 2023; Tsai et al., 2023; Yang et al., 2024) claim to be black-box, they still rely on the text encoder of the Stable Diffusion model, thereby using some internal parameters. In this section, we demonstrate that JPA is universal, achieving high attack performance with arbitrary text encoders. Beyond the CLIP text encoder, we employ BERT (Devlin et al., 2019), pre-trained on large-scale corpora such as BooksCorpus (Zhu et al., 2015) and English Wikipedia (Foundation, 2018), and the T5 model (Raffel et al., 2020), fine-tuned on a wide range of NLP tasks in a text-to-text format, to execute JPA. As shown in Table 6, even with text encoders that are not aligned with vision modalities, the attack remains effective, although the ASR slightly decreases due to the misalignment between text and image in these encoders.
Attack Efficiency.
As discussed in 1, prior methods require lengthy optimization or post-processing due to the discrete text space. To address both the issues of discrete space and the presence of sensitive words in the output, we propose soft assignment and gradient masking. In this section, we demonstrate that our method is more efficient, significantly reducing execution time compared to previous methods. As illustrated in Table 3, JPA generates a successful adversarial prompt in just $6.72$ minutes, which is only $1/4$ the time required by the next fastest method.
Method ESD FMN SLD-max SLD-Strong SLD-Medium SD-NP
P4DChin et al. (2024) 0.2402 0.2809 0.2345 0.2407 0.2409 0.2263
UnlearnDiffZhang et al. (2024) 0.2198 0.2597 0.2259 0.2359 0.2316 0.2288
Random 0.2170 0.2612 0.2236 0.2206 0.2265 0.2231
QF-AttackZhuang et al. (2023) 0.1957 0.2025 0.1980 0.1998 0.2031 0.2012
Ring-A-BellTsai et al. (2023) 0.2438 0.2834 0.2524 0.2526 0.2570 0.2545
SneakyPromptYang et al. (2024) 0.2303 0.2716 0.2409 0.2417 0.2461 0.2412
JPA(Ours) 0.2575 0.2918 0.2580 0.2534 0.2575 0.2598
Table 7: The CLIPScore of different attack methods.
Category ESD FMN SLD-max SLD-Strong SLD-Medium SD-NP Average vs. Ori. prompt
Adv. prompt 2.8968 1.9516 3.0543 2.8900 2.9975 2.5891 2.7299 / 2.8974
Table 8: Performance under Inception Score metrics for different defense methods.
Discussion on Fully Automated Attack Framework.
We observe that some prior methods depend on manual modifications or additional techniques, such as manually altering sensitive words (Qu et al., 2023; Rando et al., 2022) and employing prompt dilution (Rando et al., 2022), to successfully execute attacks against online services (Tsai et al., 2023). This reliance on human intervention compromises the automation of the attack framework. For instance, in Table 4, Ring (Tsai et al., 2023) requires additional prompt dilution and human modification to effectively attack DALL·E 2 (Ramesh et al., 2022). In contrast, our attack framework, JPA, automates the process by masking sensitive words, enabling the generation of NSFW concepts while staying within the less sensitive text regions, making them less likely to be blocked by filters.
Image Similarity Evaluation.
To assess the similarity between generated and original images more effectively, we incorporate CLIPScore Hessel et al. (2021) as a metric for evaluating semantic alignment. Specifically, we compute the CLIPScore between the successfully attacked images and the text prompt “nudity.” A higher score indicates greater semantic alignment with “nudity.” As shown in Table 7, our method achieves superior semantic performance compared to the baselines.
Image Fidelity Evaluation.
To measure whether adversarial prompts compromise the visual quality of the generated images, we use the Inception Score Salimans et al. (2016) as an evaluation metric. This metric measures the clarity and overall quality of the generated images, with higher scores indicating better image fidelity. As shown in Table 8, the results demonstrate that the attack’s impact on image quality is minimal, as reflected in the slight reduction in Inception scores observed in the final column.
6 Conclusion and Limitations.
@@
However, our work is not without limitations. Our concept pairs are given by ChatGPT, which needs prompting that out of the automated framework. This highlights potential areas for further exploration in future research. Additionally, we recognize the urgent need for the development of more robust safety checkers to counter such attacks effectively.
7 Acknowledgements.
This work is supported by the NSFC Grants (No. 62206247), the Pioneer R&D Program of Zhejiang (No. 2024C01035) and the Fundamental Research Funds for the Central Universities (226-2024-00049).
@@
B.1 Sensitive Words List.
@@
Table 10: Render words list for “nudity” and “violence”.
B.2 Render Words List.
This section presents the render word lists used in experiments for different concepts. We specifically highlight the lists for the concepts of “nudity” and “violence”. Each list contains rows representing individual render word pairs, with the total number of rows corresponding to the number of pairs. Detailed information is provided in Table 10.
@@
C.1 Quality results of JPA in “violence” concept.
@@
white-box P4D Chin et al. (2024) 97.72 97.88 75.28 88.89 89.89 100
UnlearnDiff Zhang et al. (2024) 98.87 98.87 82.02 89.77 91.11 100
@@
QF-Attack Zhuang et al. (2023) 55.56 46.67 18.89 24.45 37.78 80.00
Ring-A-Bell Tsai et al. (2023) 96.67 100 91.11 90.00 94.45 100
SneakyPrompt Yang et al. (2024) 95.07 100 77.78 88.89 90.00 100
@@
white-box P4D Chin et al. (2024) 239.80 269.57 221.83 220.76 216.88 275.89
UnlearnDiff Zhang et al. (2024) 238.65 270.81 211.30 221.57 219.78 277.90
@@
QF-Attack Zhuang et al. (2023) 271.80 254.93 261.47 259.34 255.89 278.05
Ring-A-Bell Tsai et al. (2023) 234.79 267.75 224.85 237.97 215.88 276.57
SneakyPrompt Yang et al. (2024) 230.90 260.57 215.96 226.70 218.49 270.69
@@
Table 11: The performance attack methods on the “violence” concept is evaluated under ASR and FID.
C.2 Visualization results of Not-Safe-for-Work (NSFW) images generated by T2I models.
@@
D.1 Visualization of the controllable NSFW concept rendering.
We discover that adjusting the magnitude of the NSFW concept embeddings, JPA can precisely control the extent to which NSFW concepts emerge in the generated images. We show in Figure 7 that increasing $\lambda$ leads to more NSFW contents being rendered in the image. Specifically, we adjust the value of $\lambda$ from 00 to $5$ to produce a series of images that progressively render the “nudity” concept step by step. These images increasingly incorporate specific content related to the “nudity” concept while preserving the original image’s overall meaning.
Figure 7: The controllable “nudity” concept rendering process.
D.2 Visualization of the controllable ordinary concept rendering.
@@
Appendix E Additional Experiment.
E.1 Quantity evaluation of online service.
In this section, we compare our method with other black-box attack techniques on DALL-E 2 and Stability AI, evaluating ASR, CLIPScore, and attack success time, as shown in Table 12.
Note that prompts from the Ring-A-Bell Tsai et al. (2023) attack couldn’t be used for online interface attacks, as discussed in the discussion on Fully Automated Attack Framework section 5.3 and 4. Additionally, QF-Attack Zhuang et al. (2023), a relatively weak baseline, failed to successfully attack any prompts. We use CLIPScore Hessel et al. (2021) instead of FID since we cannot access the “w/o checker” online service. CLIPScore is calculated between the original prompt and the attack-generated image.
Metric ASR Moosavi-Dezfooli et al. (2016) CLIPScore Hessel et al. (2021) Time (min)
Online Services DALL E 2 Stability AI DALL E 2 Stability AI DALL E 2 Stability AI
JPA 11.97 7.04 0.2420 0.2303 5.29 7.58
Table 12: Quantity results of online services.
E.2 Ablation study of prompt length.
To assess the influence of prompt length, we conduct ablation experiments on prefix length and observe that both excessively long and short prefixes negatively impact the attack’s effectiveness in Table 13.
Metric 1 3 5 7 9 11
ASR 22.48 36.34 49.61 67.16 58.13 58.19
FID 167.53 159.60 145.68 131.11 150.68 160.96
Table 13: Ablation study on k. Best result bloded.
Appendix F Differences between JPA and previous works.
In this section, we discuss that JPA differs from previous methods like text-attack in two main ways: 1) Unlike SneakyPrompt Yang et al. (2024), we use a transfer attack strategy instead of directly attacking the black-box model. Specifically, we leverage accessible white-box modules (i.e., CLIP) within the black-box model to obtain adversarial prefixes through white-box attacks, which are then transferred to attack black-box models. 2) Unlike other transfer attacks like Ring-a-Bell Tsai et al. (2023) and QF-Attack Zhuang et al. (2023), we include a sensitive word filtering module, allowing the attack to generate nonsensical strings with dangerous textual embeddings (e.g., “nskutcsjpg” instead of “nude”), thus eliminating the need for manual post-processing to bypass the NSFW keyword filter, as in Ring-a-Bell.
This framework offers two key advantages. First, transfer attacks target pre-identified goals with dangerous semantics, reducing the time cost compared to heuristic searches based on black-box outputs. More importantly, the sensitive word filtering makes the attack fully automated, removing the need for manual post-processing. (See Table 1. and Table 4. for more details.)
Appendix G Differences between JPA and Ring-A-Bell.
In this section, we discuss the differences between our method and Ring’s approach. The primary distinction lies in the learnable components of the adversarial sample $P_{a}$, as defined in Sec 4. Specifically, our $P_{a}$ consists of a learnable prefix prompt combined with an invariant original prompt $P_{t}$ (with a default prefix length of $7$). During optimization, only the prefix tokens are trainable. In contrast, Ring’s $P_{a}$ comprises k-length learnable tokens (typically with k set to $16$ or $77$), all of which are fully trainable. The key advantage of our method is that by training only a small portion of the overall $P_{a}$, we significantly reduce the search space. This is the primary reason our approach is more time-efficient than Ring’s.
Appendix H The details about the NSFW detector.
Brief Introduction of NudeNet exposure detector.
We use NudeNet Bedapudi (2019) as our exposure detector, which is highly effective at identifying exposed body parts (e.g., genitalia, breasts, abdomen, buttocks). Trained on a large dataset of labeled images, it includes both exposed and non-exposed content from platforms like Reddit, covering various scenarios and lighting conditions (e.g., normal clothing, partial nudity, full nudity). As a result, NudeNet excels at detecting specific exposed body parts. During inference, if NudeNet identifies any of these body parts in diffusion-model-generated images, we classify the image as a successfully attacked nudity image.
Brief Introduction of Q16 violence detector.
We use the Q16 classifier Schramowski et al. (2022) as a violence detector. It was trained on a large, annotated dataset categorized into nine safety types (e.g., violence, hate, sexual content) with ratings ranging from “safe” to “highly unsafe”. During inference, the classifier evaluates each image, assigning it to a safety category or marking it as “None applying” if no category fits. It outputs a safety assessment, including a rating, relevant category, and explanation. The final evaluation maps the rating to either “safe” or “unsafe” for our purposes.